Stats homework

A final exam in Math 10 has a mean of 73 and a standard deviation of 7.8. If 24 students are randomly selected, find the probability that the mean of their testscores is less than 76.

a. 0.7856 b. 0.5277 c. 0.8413 d. 0.9699
Answer #1

μ=73

σ=7.8

n=24

P( xbar < 76 ) = P( (xbar-μ)/(σ/√n) < (76 - 73)/(7.8/√24) ) = P( z < 1.88 ) = 1 - .0301 = .9699
Answer #2
d. .9699

answered by: Elimination method PLEASE HELP
Answer #3

μ=73

σ=7.8

n=24

P( xbar < 76 ) = P( (xbar-μ)/(σ/√n) < (76 - 73)/(7.8/√24) ) = P( z < 1.88 ) = 1 - .0301 = .9699

answered by: not too bright
Answer #4
d. .9699

answered by: Cyndy